# Rethinking Efficient and Effective Point-based Networks for Event Camera Classification and Regression: EventMamba

### v1

@@ III-B1 $LocalFE$ Extractor
The input dimension $PN_{i}$ is $[N,3+D]$, and the centroid dimension $PS_{i}$ is $[N^{{}^{\prime}},3+D]$ and the group dimension $PG_{i}$ is $[N^{{}^{\prime}},K,3+2*D]$. $K$ represents the nearest $K$ points of centroid, $D$ is the feature dimension of the points of the current stage, and 3 is the most original $(X,Y,T)$ coordinate value.
@@
III-B2 Temporal Aggregation
Following the processing of points in a group by $LocalFE$ extractor, the resulting features require integration via the aggregation module. This module is tailored to necessitate precise temporal details. Consequently, we introduce a temporal aggregation module.
@@ III-B2 Temporal Aggregation
III-B3 $GlobalFE$ Extractor
Unlike previous work[12], employing Bi-LSTM for explicit temporal information extraction nearly doubles the number of parameters, despite achieving amazing results. We incorporate the explicit temporal feature extraction module into the hierarchical structure, opting for the faster and lighter Mamba block instead. The topology of the block is shown in the green background of Figure 3. This block mainly integrates the sixth version of the State Space Model (SSM), which is well able to parallel and focus on long time series of information, such as temporal correlations between 512 event groups. The SSM extracts explicit temporal features that can be expressed as follows:
@@ III-B3 $GlobalFE$ Extractor
III-C Regreesor and loss function
@@ III-D Overall Architecture
10:     Local Extractor: Get $\mathcal{S}\in[B,T^{\prime}_{i},K,D_{i}]$
@@
12:     Temporal Aggregate: Get $\mathcal{SA}\in[B,T^{\prime}_{i},D_{i}]$
13:     $\mathcal{SA}=TAgg(\mathcal{S})$;
14:     Global Extractor: Get $PN\in[B,N_{i},D_{i}]$
15:     $PN=GlobalFE(F_{\text{aggre}})$;
@@
21:  Get 6-DOFs pose and eye coordinate $(\hat{x},\hat{y})$
IV EXPERIMENTS
In this section, we train and test the performance of EventMamba on the server platform. The experimental evaluation encompasses assessing the model’s accuracy, mean square error (MSE) performance on the datasets, analyzing the model’s parameter size, and the number of floating-point operations (FLOPs). The platform configuration comprises the following: CPU: AMD 7950x, GPU: RTX 4090, and Memory: 32GB.
IV-A Datasets
Figure 5: Event-based action recognition, CPR and eye tracking datasets visualization. Event cameras can effectively capture moving silhouettes, ignoring static objects.
We evaluate EventMamba on five popular action recognition datasets, the IJRR Camera Pose Relocalization dataset, and the SEET eye tracking dataset for classification and regression, respectively. The resolution, classes, and other information of datasets are shown in Table I.
@@ IV-A Datasets
Dataset Classes Sensor Resolution Avg.length(s) Train Samples Test Samples Sliding window Overlap
Daily DVS [45] 12 DAVIS128 128x128 3 2924 731 1.5s 0.5s
DVS128 Gesture [16] 10/11 DAVIS128 128x128 6.52 26796 6959 0.5s 0.25s
DVS Action [13] 10 DAVIS346 346x260 5 932 233 0.5s 0.25s
HMDB51-DVS [17] 51 DAVIS240 320x240 8 48916 12230 0.5s 0.25s
UCF101-DVS [17] 101 DAVIS240 320x240 6.6 110050 27531 1s 0.5s
IJRR [46] 6 DOFs DAVIS240 320x240 60 - - 5 ms 0s
SEET [37] 2 DAVIS240 240x180 20 31861 3973 4.4 ms 0s
IV-B Implement Details
IV-B1 Pre-processing
The temporal configuration for sliding windows spans intervals of 4.4 ms, 5 ms, 0.5 s, 1 s, and 1.5 s. In the specific context of the DVS128 Gesture dataset and the DVS Action dataset, the temporal overlap between neighboring windows is established at 0.25 s, while in other datasets, this parameter is set to 0.5 s. Given the prevalence of considerable noise within the DVS Action dataset during recording, we implemented a denoising method to minimize the likelihood of sampling noise points. The sliding windows of the regressed datasets are all consistent with the frequency of the labels.
IV-B2 Training Hyperparameter
@@ IV-B3 Metrics
IV-B4 Network Structure
All models use a three-stage hierarchy structure, and the number of centroids for each stage is halved, so that if the input is 1024 points, the corresponding list of centroids is [512, 256, 128]. On the other hand, as the stage increases the features of the point cloud increase, the feature dimensions are the following three lists, [32, 64, 128], [64, 128, 256], [128,256,512]. The member number of the KNN grouping is 24. Both the classifier and regressor are fully connected layers with one hidden layer of 256. The number of parameters of the model is directly related to the feature dimension of the features, and the computational FLOPs are directly related to the number of points.
@@ IV-B4 Network Structure
EventMamba 0.101 0.193 0.992
IV-C Results of Action Recognition
IV-C1 DVS128 Gesture
Table II provides a comprehensive overview of the performance and network size metrics for the DVS128 Gesture. The highest recorded accuracy, an impressive 99.6%, was reported in a frame-based approach [14]. However, it’s important to note that this method imposes substantial memory and computational demands, with a footprint approximately 120 times larger than our model. In contrast, EventMamba employs temporal feature extraction modules, resulting in a parameter size of a mere 0.101 M and a total number of FLOPs of 0.193 GFLOPs. Remarkably, this efficient model still achieves a commendable accuracy of 99.2% on the DVS128 Gesture.
@@ IV-C1 DVS128 Gesture
EventMamba 0.246 0.785 0.991
IV-C2 Daily DVS
EventMamba outperformed all other methods, achieving the same impressive accuracy of 99.1% with TTPOINT, as presented in Table III. In stark contrast, the frame-based approach, despite employing nearly 80 times more parameters, yielded a notably lower accuracy. The Event Cloud within this dataset exhibited exceptional recording quality, minimal noise interference, and distinctly discernible actions. Leveraging the Point Cloud sampling method, we are able to precisely capture the intricate motion characteristics of these actions.
@@ IV-C2 Daily DVS
Method Param.($x10^{6}$) GFLOPs Acc
@@
Method Param.($x10^{6}$) GFLOPs Acc
@@
EventMamba 0.246 0.785 0.864
TABLE VI: Model’s accuracy and complexity on UCF101-DVS.
Name Param.($x10^{6}$) GFLOPs Acc
@@
EventMamba 0.834 3.038 0.979
IV-C3 DVS Action
@@ IV-C4 HMDB51-DVS
IV-C5 UCF101-DVS
For the larger 101-class dataset, EventMamba still performs strongly, reaching an impressive 97.1%, far exceeding all previous methods. It is also very close to the 99.6% of the SOTA method [66] of UCF101 before synthesis.
@@ IV-C5 UCF101-DVS
EVTC and Event-LSTM utilize the huge model and computational resources still only approach 80% accuracy. This is a great demonstration of the superiority of fine-grained temporal information and explicit temporal features.
@@
In summary, our study employed the EventMamba model to conduct an all-encompassing evaluation of action recognition datasets, ranging from small to large scales. Our analysis demonstrated that EventMamba possesses a remarkable ability to generalize and achieve outstanding performance across datasets of varying sizes. However, after capturing explicit event temporal features, the model’s representational ability is greatly enhanced and there is a greater risk of overfitting on small datasets, such as DVS Action.
IV-D Results of CPR Regression Dataset
TABLE VII: Camera Pose Relocalization results.
Parameter 12.43M 22.35M 22.34M 16.05M 135.25M 12.63M 0.774M 0.246M
FLOPs 1.584G 3.679G 7.359G 1.822G 15.623G 1.998G 0.459G 0.392G
TABLE VIII: Data preprocessing time between two methods.
Stacking event 0.38 ms 0.53 ms 0.37 ms 0.45 ms 0.48 ms 0.53 ms 0.457 ms
Based on the findings presented in Table VII, it is apparent that EventMamba surpasses other models concerning both rotation and translation errors across almost all sequences. Notably, EventMamba achieves these impressive results despite utilizing significantly fewer model parameters and FLOPs compared to the frame-based approach. Compared to the SOTA point-based CPR network PEPNet, the computational resource consumption has also decreased, and the performance has shown a nearly 8% improvement. This underscores the pivotal role of explicit temporal feature extraction, exemplified by Mamba, in seamlessly integrating hierarchical structures. Moreover, EventMamba not only exhibits a remarkable 43% improvement in the average error compared to the SOTA frame-based CNN-LSTM method but also attains superior results across nearly all sequences, surpassing the previous SOTA outcomes. In addressing the more intricate and challenging hdr_poster sequences, while the frame-based approach relies on a denoising network to yield improved results [67], EventMamba excels by achieving remarkable performance without any additional processing. This observation strongly implies that EventMamba’s Point Cloud approach exhibits greater robustness compared to the frame-based method, highlighting its inherent superiority in handling complex scenarios.
We also present in Table VIII the time consumed for point-based and frame-based preprocessing. Point Clouds exhibit a distinct superiority over other frame-based methods in time consumption for preprocessing (sampling for point cloud and stacking events for event frame). We tested the runtime of two preprocessing methods across six sequences on our server. The frame-based methods with the addition of preprocessing time will have a significant performance penalty.
IV-E Results of Eye Tracking Dataset
Additionally, we conducted experiments with event-based eye tracking for the regression task. In contrast to the CPR task, which necessitates regression of translation and rotation in the world’s coordinate system, the eye tracking task solely requires regression of the pupil’s pixel coordinates within the camera coordinate system. To compare EventMamba’s performance more extensively, we use several representative Point Cloud networks [41, 10, 11, 12] for benchmarking.
Table IX illustrates that the optimal outcomes are highlighted in bold, whereas the second-best solutions are denoted in the underline. In comparison to classic point cloud networks, such as PointNet, PointNet++ and $\text{PointMLP}_{\text{elite}}$, EventMamba stands out by significantly enhancing accuracy and reducing distance error while maintaining few computational resources. On the other hand, CNN, ConvLSTM, and CB-ConvLSTM eye tracking frame-based methods [37] are selected for comparison. EventMamba exhibits a remarkable improvement in tracking accuracy, with a 2.4% increase in the $p_{3}$ and a 5% rise in the MSE error, while maintaining a computational cost comparable to the aforementioned methods. Besides, compared to these frame-based methods, the parameters and FLOPs in FAPNet remain constant regardless of the input camera’s resolution. That enables EventMamba to apply in different resolution edge devices.
TABLE IX: Event-based eye tracking results.
@@ IV-E Results of Eye Tracking Dataset
EventMamba 180 x 240 P 0.25 392 0.944 0.992 0.999 1.48
@@
IV-F Abation Study
IV-F1 Temporal modules
@@
To demonstrate the effectiveness of temporal modules in the CPR regression task, we replaced it with spatial aggregation and obtained average translation and rotation MSE errors of 0.018m and 1.239 in the CPR task, which is a 33.1% decrease compared to the results in the Table VII. This table also illustrates that the module responsible for extracting temporal features consumes nearly negligible computational resources, yet yields significantly enhanced performance results. This underscores the critical importance of temporal features in event-based tasks. Temporal information not only conveys the varying actions across different moments in action recognition but also delineates the spatial position of the target within the coordinate system across different temporal axis.
IV-F2 Point number

### v2

@@ 3.2.1 $LocalFE$ Extractor
The input dimension $PN_{i}$ is $[T,3+D]$, and the centroid dimension $PS_{i}$ is $[T^{{}^{\prime}},3+D]$ and the group dimension $PG_{i}$ is $[T^{{}^{\prime}},K,3+2*D]$. $K$ represents the nearest $K$ points of centroid, $D$ is the feature dimension of the points of the current stage, and 3 is the most original $(X,Y,TS)$ coordinate value. $2*D$ is to concatenate the features of the point with the features of the centroid.
@@
3.2.2 Temporal Aggregation
Following the processing of points in a group by $LocalFE$ extractor, the resulting features require integration via the aggregation module. This module is tailored to necessitate precise temporal details within a group. Consequently, we introduce a temporal aggregation module.
@@ 3.2.2 Temporal Aggregation
3.2.3 $GlobalFE$ Extractor
Unlike previous work[8], employing Bi-LSTM for explicit temporal information extraction nearly doubles the number of parameters, despite achieving amazing results. We incorporate the explicit temporal feature extraction module into the hierarchical structure, opting for the faster and stronger Mamba block instead [49]. The topology of the block is shown in the green background of Figure 3. This block mainly integrates the sixth version of the State Space Model (SSM), which is well able to parallel and focus on long time series of information, such as temporal correlations between 512 event groups. The SSM extracts explicit temporal features that can be expressed as follows:
@@ 3.2.3 $GlobalFE$ Extractor
3.3 Regreesor and Loss Function
@@ 3.4 Overall Architecture
10:     Local Extractor: Get $\mathcal{S}\in[B,T_{i-1},K,D_{i-1}]$
@@
12:     Temporal Aggregate: Get $\mathcal{SA}\in[B,T_{i},D_{i}]$
13:     $\mathcal{SA}=TAG(\mathcal{S})$;
14:     Global Extractor: Get $PN\in[B,T_{i},D_{i}]$
15:     $PN=GlobalFE(\mathcal{SA})$;
@@
21:  Get 6-DOFs pose and eye coordinate
4 EXPERIMENTS
In this section, we train and test the performance of EventMamba on the server platform. The experimental evaluation encompasses assessing the model’s accuracy, mean square error (MSE) and pixel accuracy performance on the datasets, analyzing the model’s parameter size, inference time, and the number of floating-point operations (FLOPs). The platform configuration comprises the following: CPU: AMD 7950x, GPU: RTX 4090, and Memory: 32GB.
4.1 Datasets
Figure 5: Event-based action recognition, CPR and eye tracking datasets visualization. Event cameras can effectively capture moving objects, ignoring static environment.
We evaluate EventMamba on six popular action recognition datasets, the IJRR Camera Pose Relocalization dataset, and the SEET eye tracking dataset for classification and regression. The resolution, classes, and other information of datasets are shown in Table I.
@@ 4.1 Datasets
Dataset Classes Sensor Resolution Avg.length(s) Train Samples Test Samples Point Number Sliding window Overlap
Daily DVS [50] 12 DAVIS128 128x128 3 2924 731 2048 1.5s 0.5s
DVS128 Gesture [15] 10/11 DAVIS128 128x128 6.52 26796 6959 1024 0.5s 0.25s
DVS Action [18] 10 DAVIS346 346x260 5 932 233 1024 0.5s 0.25s
HMDB51-DVS [16] 51 DAVIS240 320x240 8 48916 12230 2048 0.5s 0.25s
UCF101-DVS [16] 101 DAVIS240 320x240 6.6 110050 27531 2048 1s 0.5s
[…]
IJRR [52] 6 DOFs DAVIS240 320x240 60 - - 1024 5 ms 0s
SEET [40] 2 DAVIS240 240x180 20 31861 3973 1024 4.4 ms 0s
4.2 Implement Details
4.2.1 Pre-processing
[…]
4.2.2 Training Hyperparameter
@@ 4.2.3 Evaluation Metrics
4.2.4 Network Structure
All models use a three-stage hierarchy structure, and the number of centroids for each stage is halved so that if the input is 1024 points, the corresponding list of centroids is [512, 256, 128]. On the other hand, as the stage increases, the features of the Point Cloud increase; the feature dimensions are the following three lists: [32, 64, 128], [64, 128, 256], [128, 256, 512]. The member number of the KNN grouping is 24. Both the classifier and regressor are fully connected layers with one hidden layer of 256. The eye-tracking task has an additional sigmoid layer in the final output layer. The number of parameters of the model is directly related to the feature dimension of the features, and the computational FLOPs are directly related to the number of points discovered in Table XI.
@@ 4.2.4 Network Structure
EventMamba 0.29 0.219 0.992
4.3 Results of Action Recognition
4.3.1 DVS128 Gesture
Table II provides a comprehensive overview of the performance and network size metrics for the DVS128 Gesture. The highest recorded accuracy, an impressive 99.6%, was reported in a frame-based approach [13]. However, it’s important to note that this method imposes substantial memory and computational demands, with a footprint approximately 42 times larger than our model. In contrast, EventMamba employs temporal feature extraction modules, resulting in a parameter size of a mere 0.29 M and a total number of FLOPs of 0.219 GFLOPs. Remarkably, this efficient model still achieves a commendable accuracy of 99.2% on the DVS128 Gesture.
@@ 4.3.1 DVS128 Gesture
EventMamba 0.905 0.953 0.991
4.3.2 Daily DVS
EventMamba outperformed all other methods, achieving the same impressive accuracy of 99.1% with TTPOINT, as presented in Table III. In stark contrast, the frame-based approach, despite employing nearly 54 times more parameters, yielded a notably lower accuracy. The Event Cloud within this dataset exhibited exceptional recording quality, minimal noise interference, and distinctly discernible actions. Leveraging the Point Cloud sampling method, we are able to precisely capture the intricate motion characteristics of these actions. On the other hand, the model fitting ability becomes stronger, and we increase the point number to avoid the overfitting phenomenon.
@@ 4.3.2 Daily DVS
Method Param.($\text{x}10^{6}$) GFLOPs Acc
@@
Method Param.($\text{x}10^{6}$) GFLOPs Acc
@@
EventMamba 0.916 0.953 0.864
TABLE VI: Camera Pose Relocalization Results.
Parameter 12.43M 22.35M 22.34M 16.05M 135.25M 12.63M 0.774M 0.904M
FLOPs 1.584G 3.679G 7.359G 1.822G 15.623G 1.998G 0.459G 0.476G
Inference time 5.46 ms 6.46ms 12.46 ms 9.95ms 5.25ms - 6.8ms 5.5ms
TABLE VII: Data Preprocessing Time between Two Methods.
Stacking event 0.38 ms 0.53 ms 0.37 ms 0.45 ms 0.48 ms 0.53 ms 0.457 ms
TABLE VIII: Model’s Accuracy and Complexity on UCF101-DVS.
Name Param.($\text{x}10^{6}$) GFLOPs Acc
@@
EventMamba 3.28 3.722 0.979
[…]
Method Param($\text{x}10^{6}$) GFLOPs Acc
HMAX SNN [71] - - 32.7
Motion-based SNN [50] - - 47.3
EV-ACT [51] 21.3 29 58.5
EventMamba 0.905 0.953 59.4
4.3.3 DVS Action
@@ 4.3.4 HMDB51-DVS
4.3.5 UCF101-DVS
@@
For the larger 101-class dataset, EventMamba still performs strongly, reaching an impressive 97.1%, far exceeding all previous methods as shown in Table VIII. It is also very close to the 99.6% of the SOTA method [72] of UCF101 before synthesis. EVTC and Event-LSTM utilize the huge model and computational resources still only approach 80% accuracy. This is a great demonstration of the superiority of fine-grained temporal information and explicit temporal features.
[…]
In more challenging environments, EventMamba demonstrates superior performance over EV-ACT, a method based on multi-representation fusion, while consuming only 15% of the hardware resources demonstrated in Table IX. This underscores the compelling evidence supporting the efficacy of Point Cloud representations and our meticulously designed spatio-temporal structure.
@@
In summary, our study employed the EventMamba model to conduct an all-encompassing evaluation of action recognition datasets, ranging from small to large scales. Our analysis demonstrated that EventMamba possesses a remarkable ability to generalize and achieve outstanding performance across datasets of varying sizes. However, after capturing explicit event temporal features, the model’s representational ability is greatly enhanced and there is a greater risk of overfitting on small datasets, such as DVS Action and Daily DVS.
TABLE X: Event-based Eye Tracking Results.
@@
EventMamba 180 x 240 P 0.903 476 0.944 0.992 0.999 1.48
$\text{EventMamba}_{\text{lite}}$ 180 x 240 P 0.27 62.6 0.936 0.986 0.998 1.45
@@
TABLE XI: Ablation Study on HMDB51-DVS Dataset.
Point Number Dimension [16, 32, 64] [32, 64, 128] [64, 128, 256] 256 w/o Mamba 256 w LSTM
512 points (48.2 %, 0.102 M, 17 M) (58.6 %, 0.28 M, 63 M) (67.9 %, 0.92 M, 238 M) ( 43.9 %, 0.26 M, 191 M) ( 58.3 %, 0.784 M, 229M)
1024 points (52.7 %, 0.102 M, 35 M) (67.8 %, 0.28 M, 125 M) (76.6 %, 0.92 M, 476 M) ( 52.7 %, 0.26 M, 383 M) ( 61.4 %, 0.784 M, 459M)
2048 points (59.2 %, 0.102 M, 71 M) (74.3 %, 0.28 M, 252 M) (82.1 %, 0.92 M, 953 M) ( 57.5 %, 0.26 M, 767 M) ( 70.8 %, 0.784 M, 919M)
TABLE XII: Abation of Temporal Modules on IJRR Dataset.
Network w/o Mamba w LSTM EventMamba
Parameter 0.245M 0.774M 0.904M
FLOPs 0.392G 0.459G 0.476G
shapes_rotation 0.007m,1.481 0.005m,1.372 0.004m,1.091
box_translation 0.019m,1.237 0.017m,0.845 0.016m,0.810
shapes_translation 0.015m,0.827 0.011m,0.582 0.010m,0.600
dynamic_6dof 0.018m,1.154 0.015m,1.045 0.014m,0.911
hdr_poster 0.029m,1.744 0.016m,0.991 0.015m,0.985
poster_translation 0.021m,0.991 0.012m,0.588 0.013m,0.587
Average 0.018m,1.239 0.013m,0.904 (+27.3%) 0.012m,0.831(+33.1%)
4.4 Results of CPR Regression Dataset
Based on the findings presented in Table VI, it is apparent that EventMamba surpasses other models concerning both rotation and translation errors across almost all sequences. Notably, EventMamba achieves these impressive results despite utilizing significantly fewer model parameters and FLOPs compared to the frame-based approach. Compared to the SOTA point-based CPR network PEPNet, the computational resource consumption preserves consistency, inference speed increased by 20%, and the performance has shown a nearly 8% improvement. This underscores the pivotal role of explicit temporal feature extraction, exemplified by Mamba, in seamlessly integrating hierarchical structures. Moreover, EventMamba not only exhibits a remarkable 43% improvement in the average error compared to the SOTA frame-based CNN-LSTM method but also attains superior results across nearly all sequences, surpassing the previous SOTA outcomes. In addressing the more intricate and challenging hdr_poster sequences, while the frame-based approach relies on a denoising network to yield improved results [33], EventMamba excels by achieving remarkable performance without any additional processing. This observation strongly implies that EventMamba’s Point Cloud approach exhibits greater robustness compared to the frame-based method, highlighting its inherent superiority in handling complex scenarios.
We also present in Table VII the time consumed for point-based and frame-based preprocessing. Point Clouds exhibit a distinct superiority over other frame-based methods in time consumption for preprocessing (sampling for Point Cloud and stacking events for event frame). We tested the runtime of two preprocessing methods across six sequences on our server. The frame-based methods with the addition of preprocessing time will have a significant performance penalty.
4.5 Results of Eye Tracking Dataset
Additionally, we conducted experiments with event-based eye tracking for the regression task. In contrast to the CPR task, which necessitates regression of translation and rotation in the world’s coordinate system, the eye tracking task solely requires regression of the pupil’s pixel coordinates within the camera coordinate system. To compare EventMamba’s performance more extensively, we use several representative Point Cloud networks [45, 10, 11, 8] for benchmarking. Table X illustrates that the optimal outcomes are highlighted in bold, whereas the second-best solutions are denoted in the underline. In comparison to classic Point Cloud networks, such as PointNet, PointNet++ and $\text{PointMLP}_{\text{elite}}$, EventMamba stands out by significantly enhancing accuracy and reducing distance error, and $\text{EventMamba}_{\text{lite}}$ also maintains high performance while consuming few computational resources. On the other hand, CNN, ConvLSTM, and CB-ConvLSTM eye-tracking frame-based methods [40] are selected for comparison. our models exhibit a remarkable improvement in tracking accuracy, with a 3% increase in the $p_{3}$ and a 6% decrease in the MSE error, while maintaining a computational cost comparable to the aforementioned methods. Besides, compared to these frame-based methods, the parameters and FLOPs in EventMamba remain constant regardless of the input camera’s resolution. That enables EventMamba to apply in different resolution edge devices.
4.6 Abation Study
4.6.1 Temporal modules
@@
To demonstrate the effectiveness of temporal modules in the CPR regression task, we replaced it with spatial aggregation and obtained average translation and rotation MSE errors of 0.018m and 1.239 in the CPR task, which is a 33.1% decrease compared to the results in the Table XII. This table also illustrates that the module responsible for extracting temporal features consumes nearly negligible computational resources yet yields significantly enhanced performance results. On the other hand, Table XI demonstrates that the Point Cloud network without a temporal module only achieves 57.5% accuracy on the HMDB51-DVS dataset. These underscore the critical importance of temporal features in event-based tasks. Temporal information not only conveys the varying actions across different moments in action recognition but also delineates the spatial position of the target within the coordinate system across different temporal axes.
4.6.2 Point Number